4. Use the Laplace transform to solve the initial value problem y + y = f(1) = -2, ost<2 13t+4, 122 y(0) = 0, y(0) = -1
Answer #1

T LCg+Ly) - Sest corde + Jest(++4206 selly)-sy(0)=y607+ Ley) = REST* + (s++4229 lei (5+10LLY)--+ = les.cº) = [ (36+4)est +ACS +1 +5(85+(?=1 SPA+B)+CS+A =1 Azl, CEO, A+B=0 B=-A=7 1884 17 = 32 st. (841)-5² Ⓡ 040 541841) 1). -